Average speed equals distance divided by time. On a journey to the Alpha Centauri star system, a space probe travels a distance of 4.12×10[sup]16[/sup] meters in 1.73×10[sup]10[/sup] seconds. (a) Find its average speed in meters per second. (b) How long did the journey take in years? Use 365.24 days per year
1
Expert's answer
2011-02-01T04:51:51-0500
a) V = S/t = 4.12×1016 [m] / 1.73×1010 [s] = 2381502.89 m/s
b) In the year there is 365.24 days, in the day there is 24 hours, in one hour - 60 minutes, in one minute - 60 seconds; thus
1.73*1010/(365.24*24*60*60) = 548.22 years.